The table below shows the values for several different components of GDP.

Value in billions of dollars
personal consumption expenditures: 265
Net foreign factor income: 4
Transfer Payments: 14
Rents: 14
Consumption of fixed capital (depreciation): 19
Statistical discrepancy: 8
Social Security contributions: 12
Interest: 14
Proprietors Income 25
Net exports: 14
Dividends 16
Compensation of employees 243
Taxes on production and imports: 10
Undistributed corporate Profits 13
Personal taxes: 18
Corporate taxes 11
Corporate profits 48
Government purchases: 54
Net private domestic investment: 25

Personal saving: 12

Imports: 7

12

7

A: What is the value of exports?

B: What is the value of National Income?

C: What is the value of GDP?

D: What is the value of personal income?
Answer #1

A.

NX=X-M

14=X-7

X=14+7

=21

B.

National Income= GDP- depreciation+ net factor income from abroad – indirect business tax

=358-19+4-10

=$333

C.

GDP by expenditure method=C+I+G+X-M

=265+25+54+14

=$358

D.

Personal income=GDP– depreciation – retained earning

=358-19-13

=$326